International Congress of Mathematicians Thursday, August 22, 2002 – No. 3

Announcement

Program Committee:

Chair: Yuri Manin, Max-Planck-Institut fur Math., Bonn John Ball, University of Oxford Franco Brezzi, University of Pavia Gerard Laumon, Univ. de Paris-Sud Laszlo Lovasz, Yale University Tetsuji Miwa, Kyoto University Peter Sarnak, Princeton University Alain-Sol Sznitman, ETH-Zentrum, Zurich Gang Tian, Massachusetts Institute of Technology Michele Vergne, Ecole Normale Superieure, Paris Wen-Tsun Wu, Academia Sinica, Beijing

Fields Medal Committee:

Chair: Yakov Sinai, Princeton University James Arthur, University of Toronto Spencer Bloch, University of Chicago Jean Bourgain, Institute for Advanced Study Helmut Hofer, ETH-Zentrum, Zurich Yasutaka Ihara, Kyoto University H. Blaine Lawson, SUNY Stony Brook Sergei Novikov, University of Maryland George Papanicolaou, Stanford University Efim Zelmanov, Yale University Nevanlinna Committee:

Chair: Michael Rabin, Harvard University Andrei Agrachev, Steklov Institute Ingrid Daubechies, Princeton University Wolfgang Hackbusch, University of Kiel Alexander Schrijver, CWI, Amsterdam

August 22, 16:00-17:00, CH17C. Lecture of the Fields Medalist: Voevodsky, Vladimir (Institute of Advanced Study, USA), The slice filtration in motivic stable homotopy. followed by Informal Seminar on K-Theory: 17:15-18:00, CH17C, Geisser, Thomas (University of Southern California, USA), Weil etale motivic cohomology. 18:00-18:45, CH17C, Kahn, Bruno, (University of Paris 7, France), Algebraic K-theory, algebraic cycles, and arithmetical geometry.

August 23, 19:30-20:30, MR6-8.

Lecture of the Nevanlinna Prize Winner: Madhu, Sudan, (MIT, USA), List decoding of error-correcting code.

Scientific Program – Changes

1. Logic Wencel, Roman, From Aug. 24, 16:45–17:00, R3058 ⇒ Aug. 22, 14:25–14:40, R3058.

2. Algebra Title changed: Aug. 22, 17:00–17:15, Artamonov, Viatchesalv A., On quantum polynomials, R5021.

7. Lie Groups and Representation Theory The chair of Eckhard Meinrenken’s lecture is changed: Zhao, Kaiming, (AMSS, CAS, China), Aug. 23, 17:15-18:00, CH17B.

8. Real and Complex Analysis

Changed (Poster to Short talk): Sukhotin, Alexander Mikhailovich, From Aug. 27, 14:45–15:00, R3051 ⇒ Aug. 24, 17:00–17:15, R5015.

9. Operator Algebras and Functional Analysis Niknam, Assadollah (Ferdowsi University of Mashad, Iran), From Aug. 27, 14:45–15:00, R3051 ⇒ Aug. 24, 14:25–14:40, R5015.

18. Mathematics Education and Popularization of Mathematics Short Talk: Aug. 24, 15:30–15:45, Michtchenko, Tatiana M., Basic geometric configurations as a principle of methodical maintenance of the school geometry course, R3058, ⇒ Poster: Aug. 24, 17:20–18:20, EH02.

ad-hoc Talks and Posters

1. Logic Short Talk: Welch, Philip (Bonn & Bristol), On Foreman & Magidors’ mutually stationary sets, Aug. 24, 16:45–17:00, R3058.

2. Algebra Short Talk: Shahabi, Mohammad Ali (Shojaei Tabriz University, Iran), Cycle structure of the group SP(4,q), Aug. 23, 15:15–15:30, R5023.Short Talk: Neelakanta, Sthanumoorthy (University of Madras, India), Some classes of root systems of generalized Kac-Moody algebras, Aug. 23, 15:30–15:45, R5023.

3. Number Theory Short Talk: Ongarbaev, Ernar Saginbekovitch (Yasawi Kazakh-Turkish International University, Kazakhstan), To the Golbach-Euler’s theorem, Aug. 23, 15:30– 15:45, CH14.

4. Differential Geometry Poster: Olszak, Zbigniew (Wroclaw University of Technology, Poland), Para-Hemitian stuctures on twisted products of Lie groups, Aug. 22, 14:45–15:45, EH02. Poster: Deszcz, Ryszard (Agricultural University of Wroclaw, Poland), Hypersurfaces in semi-Riemannian space forms satisfying some curvature conditions, Aug. 22, 14:45–15:45, EH02.

5. Topology Poster: Namdari, Mehrdad (Shahid Chamran University of Ahvaz, Iran), Algebraic properties of C_{∞}(X), Aug. 24, 14:45–15:45, EH02. Short Talk: Tserennadmid, Batkhuu (Mongolian State Pedagogical University, Mongolia), Ω-box topology, Aug. 24, 4:45–15:00, R3051.

9. Operator Algebras and Functional Analysis Short Talk: Smailov, Yesmukhanbet S. (Karaganda State University of E.A.Buketov, Kazakhstan), About embedding in space of multiplicators of trigonometric Fourier series, Aug. 23, 14:25–14:40, R5021.

10. Probability and Statistics

Short talk: Fourati, Sonia (Universite Paris 6, France), Wiener-Hopf factorisation of Cauchy processes and the Dilogarithm, Aug. 24, 18:05–18:20, R5020.

11. Partial Differential Equations

Short talk: Kim, Wan Se (Hanyang University, Korea), Multiplicity result for nonlinear parabolic equations having singular nonlinearity, Aug. 22, 17:00–17:15, R3051.

12. Ordinary Differential Equations and Dynamical Systems Short talk: Riahi, Hasna (Ecole Nationale d’Ingenieurs de Tunis, Tunisia), On the parabolic-like orbits for some sigular Hamiltonian systems, Aug. 22, 14:25–14:40, CH03.Short talk: Aulbach, Bernd (University of Augsburg, Germany), Geometric Properties of Nonautonomous Dynamical Systems, Aug. 23, 17:20–17:35, R5021. Short talk: Kakishov, Kanybek Kakishovich (Kyrgyz State National University, Kyrgyzstan), Aug. 24, 14:25–14:40, R5023.

13. Mathematical Physics

Short talk: Leng, Wilfrid S. (University of Cambridge, United Kingdom), Settled and unsettled fine-topologies with applications to the poincare problem and to the standard model of mathematical physics, Aug. 22, 17:35–17:50, R3030.

Cancellation

9. Operator Algebras and Functional Analysis Aug. 22, 14:45–15:00, Chen, Fangqi (Tianjin University, China), Global solutions for nonlinear impulsive Volterra integral equations in Banach spaces, R5015.

10. Probability and Statistics Aug. 22, 16:30–16:45, Shahumyan, Harutyun (National Academy of Sciences, Armenia), On the most informative indices selection method, R5023. Aug. 27, 13:55–14:10, Sun, Liuquan, Regression analysis for a semi-parametric model with panel data, R5020.

12. Ordinary Differential Equations and Dynamical Systems Aug. 23, 17:20–17:35, Koplatadze, Roman (A. Razmadze Mathematical Institute of the Georgian Academy of Sciences, Georgia), Comparison theorem for differential equations with several deviations, R5021.

13. Mathematical Physics

Aug. 26, 17:20–17:35, Fan, Engui, A family of completely integrable multi-Hamiltonian structure systems explicitly related to some celebrated equations, CH13.

Footloose tours – Item I

A brief of descriptions of the footloose tours on August 22 are as follows. Tour 3: Chinese Ethnic Culture Park; booked out Description: This culture park is indeed a big museum of the nationalities of China. China has 56 nationalities. In this park, houses(villages) of all the different nationalities can be found. Culture shows, such as special dances, wedding ceremonies, songs, festival celebrations of different ethnic groups, are given in the park regularly.A visit to this park will at least give you a flavour of China’s multinational cultures. Meeting: 9:00 a.m. in front of BICC Duration: aprrox. 3 hours Entrance Fee: 60 yuans Tour 10: Legend Computer Company; booked out Description: Legend is the best selling PC brand in the PRC, commanding 30% of the market share in 2001. Its innovative policies and strategic alliances with international computer vendors, continue to reaffirm its leading position in the PRC’s computer market. In fact, It is regarded by many of the international vendors as a key partner. This tour will take you to the headquarters Office of the Legend Company. Meeting: 9:00 a.m. in front of BICC Duration: aprrox. 3 hours Entrance Fee: free Tour 6-A: Yuan Ming Yuan (the old Summer Palace); booked out Description: Yuan Ming Yuan (the Garden of Perfection and Brightness) is actually three gardens. It used to be the summer palace in Qing Dynasty (please notice that it is older than the Summer Palace, the later one being one of the most visited places in Beijing now). The whole garden was destroyed by Anglo-French joint force in 1860. The whole grounds were set on fire, and treasures were plundered. Since most the buildings were wooden structures, almost nothing escaped the fires, except in on corner of the garden there still stand some broken stone pillars and collapsed arches. Meeting: 15:00 p.m. in front of BICC Duration: aprrox. 3 hours Entrance Fee: 10 Yuans for the Park, an additional 15 Yuans to see the Ruins.

Footloose tours – Item II

LOCAL TOURS LT-1: The Great Wall at Juyong Pass and the Ming Tombs Date: August 22. Description: The Great Wall stretches 6000 kilometers from East to West China, traversing five provinces and two autonomous regions. Built first in 221 B.C. and re-built several times in the following 1500 years, it has become a symbol of the ancient Chinese civilization. After the visit to the Great Wall and lunch, you will visit an emperor’s tomb of the Ming Dynasty (1380-1644). This tomb is a mausoleum of the 13th Ming emperor and is the only one that has been excavated. And, the best part of the tomb area is the Spirit Way - giant stone carvings of human beings, animals and mythical monsters standing guard. Departure time: 08:30 in front of BICC Fees: USD 38 or RMB 315 per person, lunch included Duration: 08:30- 17:00 LT-3: Summer Palace and the Lama Temple Description: Summer Palace, the largest and best preserved imperial garden. It used to be a retreat for emperors to escape the scorching summer heat in Beijing. The garden is so large that it is scarcely enough to see in one day all the halls, towers, pavilions, corridors, walkways and bridges that grace the hill slopes, lake shore and islets. In the afternoon, you will visit the Lama Temple. The temple is the most impressive and an active Buddhist temple in the city. One of the prominent features of the temple is the 23-meter-tall Maitreya Buddha. Departure time: 08:30 in front of BICC Fee: USD 26 or RMB 215 per person, lunch included. Duration: 08:30- 16:30 LT-4: Tiananmen Square, Beihai Park and the lane tour Description: Tiananmen Square is located in the center of the city, to the south of the Forbidden City. It is the largest city square in the world. At the Square there are Mao’s Mausoleum, the Museum of Chinese History and the Great Hall of the People (parliament building). Beihai Park is to the northwest of the Forbidden City. It was built on the basis of a legend and used to be a royal garden. After lunch, you will visit the backyards of Beijing, on a pedicab, across the small zig-zag lanes. Departure time: 08:30 in front of BICC Fees: USD 26 or RMB 215 per person, lunch included. Duration: 08:30-16:30

New Footloose Tour

A new footloose tour will be arranged on Sunday, August 25, 2002. The tour will take you to Nankai University in Tianjin. Nankai University is one of the top universities in China. The mathematics institute of Nankai University was founded by Prof Shiing-Shen Chern in 1985. The bus will depart from BICC at 9.00AM, it will return BICC at about 5.30PM.

There is a free postcard for each ICM2002 participant. Please go to R3055 third floor of the Convention Hall to receive the postcard by showing your name tag and ID number.

Daily News of the

International Congress of Mathematicians, August 20–28, 2002, Beijing, China Editor in Chief: Wenlin Li

Editors:

Qi Feng, Shirong Guo, Yafen Jin, Huimin Weng, Huijuan Wang, Qixiao Ye English Consultant: Joseph C.Y. Chen

Latex Support: Jinrong Wu